# QCResUNet: Joint Subject-level and Voxel-level Segmentation Quality Prediction

### v1

@@ 1 Introduction
Medical image segmentation plays an indispensable role for accurate diagnosis, monitoring, treatment planning, and population studies of diseases in modern medicine by enabling precise delineation of anatomical structures and pathological regions (Garcia-Garcia et al., 2017; Litjens et al., 2017). In particular, precise segmentation of healthy and abnormal anatomy into multiple classes using magnetic resonance imaging (MRI) is crucial in this process. Recently, deep learning-based methods have achieved state-of-the-art performance in automated segmentation tasks including brain tumor (Ronneberger et al., 2015; Kamnitsas et al., 2017; Isensee et al., 2018a, b; Baid et al., 2021) and cardiac (Tran, 2016; Khened et al., 2019; Zhou et al., 2021) MRI segmentation. However, deep neural networks are sensitive to data distribution. This makes them prone to reduced performance when applied on out-of-distribution MRI scans due to variations in acquisition protocols, contrast, image quality, etc. Therefore, quality control (QC) is necessary to thoroughly assess segmentations before they are used for clinical purposes or large-scale research studies. QC tools are required to detect severe segmentation failures on a per-case basis, pinpoint areas needing segmentation refinement at the voxel level, and provide a quality measure for downstream analyses. Previously proposed QC methods fall into four main categories: uncertainty estimation-based, generative model-based, reverse classification accuracy (RCA)-based, and regression-based.
@@
The first category of automated segmentation QC methods operates on the premise that high uncertainty is reflective of poor quality segmentations (Ng et al., 2018; Albà et al., 2018; Sander et al., 2020; Ng et al., 2020; Bai et al., 2018; Roy et al., 2018; Jungo et al., 2020; Mehta et al., 2022). Accordingly, most studies focused on developing uncertainty measures to aggregate voxel-wise uncertainty as a proxy for segmentation quality measures, such as Dice Similarity Coefficient (DSC). However, most developed proxy measures (Ng et al., 2018; Roy et al., 2018; Jungo et al., 2020) have not demonstrated a strong correlation with DSC. Moreover, non-negligible errors in uncertainty estimation were observed at the voxel level resulting in unreliable subject-level uncertainty aggregation (Jungo et al., 2020). Additionally, these methods can only be applied to deep learning-based models, and thus cannot be used to assess the quality of segmentations obtained by other methods.
@@
The second category of automated QC methods is based on the assumption that there is a relationship between image intensities and tissue labels. Along that direction,  Grady et al. (2012) proposed a Gaussian mixture model with hand-crafted features (i.e., geometric features, intensity features, gradient features, and ratio features) to characterize the segmentation quality and detect segmentation failure.  Wang et al. (2020b) proposed a variational autoencoder (VAE) to learn the latent representation of pairs of images and their ground-truth segmentations. During inference, the encoder is frozen while the decoder is refined for a given image-segmentation pair to produce a surrogate segmentation. The subject-level DSC is then computed between the query segmentation and the surrogate segmentation. The tuning of the decoder is required for each query image, which may be computationally expensive and time-consuming. Leveraging the advances in the image-to-image translation task, Li et al. (2022) proposed a generative adversarial network (GAN) to generate an informative reference image conditioned on the query segmentation mask that needs to be assessed. An auxiliary network (i.e., difference investigator) is then trained to predict image-level and pixel-level quality by taking the raw input image and the generated reference image as inputs. These approaches have only been validated in cardiac MRI segmentation QC that involves only a single modality and segmentation with regular shapes. Translating it to the more complex brain tumor segmentation QC scenario is difficult due to the presence of multiple modalities and intratumoral tissue heterogeneities, which are challenging to model.
@@
The third category of the automated segmentation QC methods is built upon the reverse classification accuracy (RCA) framework (Valindria et al., 2017). This was initially designed for whole-body multi-organ MRI segmentation QC and was later applied to cardiac MRI datasets (Robinson et al., 2019). The RCA framework involves: (i) choosing a reference dataset with known ground-truth segmentation, (ii) training a segmenter using a query image-segmentation pair, (iii) using the trained segmenter to segment the images in the reference dataset, and (iv) estimating DSC as the maximum DSC achieved by the trained segmenter in the reference dataset. Although effective for whole-body multi-organ MRI segmentation and cardiac MRI segmentation QC, RCA’s reliance on a representative reference dataset and image registration poses challenges in brain tumors. The large variability in brain tumors makes the reference dataset hard to be representative. Different from whole-body multi-organ and cardiac segmentation, which have consistent shapes and appearances, the heterogeneous appearance and phenotypes of brain tumors complicate the establishment of correspondences for the tumorous areas between different subjects. Lastly, despite predicting subject-level DSC, the RCA framework lacks segmentation error localization at the voxel level.
@@
The fourth category of the segmentation QC methods is regression-based methods that directly predict the subject-level DSC. Early attempts employed a Support Vector Machine regression in combination with hand-crafted features to detect cardiac MRI segmentation failures (Kohlberger et al., 2012; Albà et al., 2018). Instead of using hand-crafted features,  Robinson et al. (2018) proposed a convolutional neural network (CNN) regressor to automatically extract features from a large cardiac MRI segmentation dataset to predict DSC. Besides predicting DSC, Kofler et al. (2022) proposed a holistic rating to approximate how expert neuroradiologists classify high-quality and poor-quality segmentations. However, manually deriving holistic ratings by clinical experts is laborious and prone to inter-rater variability, hindering its applicability in large-scale datasets. Despite the fact that satisfactory performance was achieved in predicting subject-level segmentation metrics by these regression-based methods, voxel-level localization of segmentation error is unavailable.
@@
Although numerous efforts have been devoted to automated segmentation QC, the majority of the studies have focused on whole-body multi-organ MRI segmentation and cardiac MRI segmentation QC. Additionally, most of these approaches have not assessed out-of-distribution generalization. Importantly, there has been limited research on segmentation QC specifically for brain tumor MRI segmentation QC. The heterogeneous and complex appearances of brain tumors, which vary in locations, sizes, and shapes, contribute to the increased challenge of QC in brain tumor MRI segmentation. In addition, most prior QC approaches focused on predicting subject-level DSC (Jungo et al., 2020; Robinson et al., 2018, 2019; Valindria et al., 2017; Li et al., 2022) and do not consider the quality of the segmentation contour. However, DSC and contour measurement (e.g., normalized surface dice (NSD)) are equally important to comprehensively assess the segmentation quality (Maier-Hein et al., 2024). Furthermore, it is crucial to address reliable voxel-level tissue-specific segmentation error localization. This localization is vital not only for auditing purposes but also for radiologists to prioritize cases that require manual refinement. By incorporating this component into quality control measures, it can enhance the accuracy and effectiveness of the overall segmentation process. However, this aspect has been largely overlooked in previous studies. One exception is the work proposed by Li et al. (2022). However, this approach was only validated on cardiac segmentation QC with limited datasets. In addition, this method can only provide a binary segmentation error mask and is unable to identify voxel-level segmentation errors for different tissue classes.
@@
Figure 1: Data generation for the brain tumor segmentation QC task: (a) The histogram of the DSC distribution in the generated BraTS training dataset before and after applying the resampling strategy. (b) Visual examples of the generated segmentation dataset ranging from low-quality to high-quality.
@@
To address these limitations, we proposed a novel deep learning model, termed QCResUNet, to jointly predict segmentation-quality metrics at the subject level and localize segmentation errors across different tissue classes at the voxel level. This work extends our previous preliminary work (Qiu et al., 2023) in several ways. First, we extended the previous work to predict subject-level DSC and NSD as well as a collection of binary segmentation error maps, each corresponding to a different tissue class, to provide a more comprehensive evaluation of segmentation quality. To achieve this goal, we further proposed an attention-based segmentation error map aggregation mechanism to better delineate segmentation errors for different tissue classes. Second, we conducted rigorous evaluations to validate the generalizability of the proposed QCResUNet by performing a three-fold cross-validation on the internal dataset and evaluating the proposed method on out-of-distribution datasets with MRI scans of varying image quality and segmentations produced. Third, we further examined the generalizability of the proposed QCResUNet by evaluating its ability to assess the segmentation quality for cardiac MRI. Fourth, we extensively evaluated the performance of the proposed method in comparison with several state-of-the-art segmentation QC methods, including RCA-based (Valindria et al., 2017) and uncertainty-based methods (Jungo et al., 2020). Fifth, we performed an in-depth explainability analysis to elucidate the reasons behind the proposed method’s superior performance compared to other competing approaches.
The main contributions of this work are fourfold:
@@ 2.1 Dataset
In this study, we used three datasets for evaluation of the QC performance on brain tumor segmentation. First, we used pre-operative multimodal MRI scans with gliomas of all grades (WHO Central Nervous System grades 2-4) grades from the BraTS 2021 challenge training dataset ($n=1251$). The BraTS dataset is a heterogeneous dataset consisting of cases from 23 different sites with various levels of quality and protocols. The BraTS dataset was used for training, validation, and internal testing. Additionally, we used two datasets that are not included in the BraTS 2021 dataset (i.e., the BraTS-SSA dataset and the WUSM dataset) for external testing, allowing for an unbiased assessment of the generalizability of the proposed method. The BraTS-SSA dataset ($n=40$(Adewole et al., 2023) is an extension of the original BraTS 2021 dataset with patients from Sub-Saharan Africa, which includes lower-quality MRI scans (e.g., poor image contrast and resolution) as well as unique characteristics of gliomas (i.e., suspected higher rates of gliomatosis cerebri). The WUSM dataset ($n=175$) was obtained from the retrospective health records of the Washington University School of Medicine (WUSM), with a waiver of consent, in accordance with the Health Insurance Portability and Accountability Act, as approved by the Institutional Review Board (IRB) of WUSM (IRB no. PA18-1113). Each subject in all datasets comprised four modalities viz. pre-contrast T1-weighted (T1), T2-weighted (T2), post-contrast T1-contrast (T1c), and Fluid attenuated inversion recovery (FLAIR). In addition, multi-class tumor segmentation masks annotated by experts were also available. Segmentation masks delineated enhancing tumor (ET), non-enhancing tumor core (NCR), and edema (ED) classes. Following standard BraTS procedures, we combined the binary ET, NCR and ED segmentation masks to delineate the whole tumor (WT), tumor core (TC), and enhancing tumor. The WT mask consists of all tumor tissue classes (i.e., ET, NCR, and ED), while the TC mask comprises ET and NCR tissue classes.
@@
Scans from the BraTS training and BraTS-SSA datasets were already registered to the SRI24 anatomical atlas (Rohlfing et al., 2010), resampled to 1-mm3 isotropic resolution and skull-stripped. For consistency, raw MRI scans from WUSM were pre-processed following the same protocol using the Integrative Imaging Informatics for Cancer Research: Workflow Automation for Neuro-oncology (Chakrabarty et al., 2022) framework. Subsequently, we z-scored all the skull-stripped scans in the BraTS datasets and the WUSM dataset on a per-scan basis. Finally, scans from the entire dataset were cropped to exclude background regions and then were zero-padded to a common dimension of $160\times 192\times 160$ using the nnUNet preprocessing pipeline (Isensee et al., 2018b).
@@
In evaluating the QC performance on the cardiac segmentation task, we used the Automated Cardiac Diagnosis Challenge (ACDC) dataset (Bernard et al., 2018), which consists of 100 subjects (200 volumes). Each volume in the ACDC dataset is associated with a multi-class segmentation mask delineating the left ventricle (LV), myocardium (Myo), and right ventricle (RV). Lastly, each volume was cropped and zero-padded to a common dimension of $16\times 16\times 160$ using the nnUNet pipeline (Isensee et al., 2018b).
@@ 2.2 Segmentation Dataset Generation
As far as the brain tumor segmentation QC task is concerned, we first employed both CNN-based (i.e., nnUNet (Isensee et al., 2018b)) and Transformer-based (i.e., nnFormer (Zhou et al., 2021)) segmentation UNets and trained them independently seven times by taking different modalities as input (namely T1-only, T1c-only, T2-only, FLAIR-only, T1-FLAIR, T1c-T2, and all four modalities). Since each modality is sensitive to certain tissue types (e.g., T1c is effective in detecting enhancing tumors and FLAIR is advantageous in identifying edema), this approach enabled us to produce segmentations with varying levels of quality. Second, we sampled the segmentation along the training routines at various iterations to generate segmentation results from both fully trained and inadequately trained models. For this purpose, a small learning rate ($1~{}\times~{}10^{-6}$) was used during training to slower their convergence, allowing the segmentation to improve gradually from low quality to high quality. Third, we employed a data augmentation strategy based on image transformations to produce diverse segmentation results, termed SegGen. A series of image transformations, including rotation, translation, scaling, and elastic deformation, were randomly applied to each ground-truth segmentation three times with a probability of $0.5$ (see Table 1). Fourth, apart from the segmentations produced by the aforementioned three methods, we also created out-of-sample segmentation samples from a different segmentation method to assess the generalizability of our proposed model in the testing phase. To accomplish this, we employed the DeepMedic framework (Kamnitsas et al., 2017) to train seven segmentation models using the same diverse input modalities as we did in training the nnUNet and nnFormer. After applying the aforementioned four steps, we were able to create a set of erroneous segmentations for the cases included in the internal and external datasets. The erroneous segmentations along with the ground truth segmentations were subsequently used to estimate the DSC, NSD, and the ground truth segmentation error map (SEM) that were used to train the proposed network. The SEM was computed as the difference between a query segmentation (i.e., one of the erroneous segmentations we produced) and the corresponding ground truth. Examples of the generated segmentations can be found in Figure 1(b).
@@
Figure 2: (a) The proposed QCResUNet model is a U-shaped neural network that takes as input $m$ imaging modalities ($\{M_{1},M_{2},\cdots,M_{m}\}$) and a multi-class segmentation mask to be evaluated ($S_{query}$). It generates three outputs: the subject-level segmentation-quality metrics DSC ($\text{DSC}_{pred}$) and NSD ($\text{NSD}_{pred}$) as well as a collection of $C$ binary voxel-level segmentation error maps $\{\text{SEM}_{\text{tissue}_{c}}\}_{c=1}^{C}$, each for each tissue class. In the case of brain tumor segmentation task, which is demonstrated as an example in this figure, QCResUNet takes as input four imaging modalities ($\{M_{1},M_{2},M_{3},M_{4}\}$) and a query segmentation $S_{query}$ that delineates WT, TC, and ET. It produces subject-level DSC and NSD, along with three binary SEMs corresponding to the segmentation error masks for WT, TC, and ET tissue classes. Note that the depicted image sizes as well as the number and size of convolution filters are specific to the brain tumor segmentation tasks. Subpanel figures (b) (c), and (d) depict the Residual Block employed in the encoder of QCResUNet, the Convolutional Block used in its decoder, and the Efficient Channel Attention (ECA) used for multiclass SEM aggregation, respectively. The abbreviations in the figure: Conv3D = 3D convolutional layer, GAP = global average pooling, FC = fully connected layer, LeakyReLU = leaky rectified linear unit, and One-hot Encoding = one-hot encodes the multi-class $S_{query}$ to a collection of binary masks.
@@
Despite our efforts to sample across different convergence stages by adopting a small learning rate, we observed that the deep learning models could successfully segment the majority of the cases. As a consequence, the resulting dataset was distributed unevenly across the different levels of quality, comprising mostly of higher quality segmentation results (see Figure 1(a)). Following the resampling strategy in Robinson et al. (2018), we randomly selected a subset of samples from each bin of the DSC histogram. This ensured that the number of segmentations from each bin was equal to the lowest count-per-bin value ($n_{s}$) across the distribution for all the segmentation datasets (refer to Figure 1(a)). For this purpose, we divided DSC values into 10 evenly spanned bins ranging from 0 to 1. We sampled cases from these bins such that we ensured that all bins contained an equal number of cases. For the training set, we did not manually select the DSC samples. Instead, at each training iteration, we randomly sampled segmentations that were evenly distributed across 10 bins to avoid imbalanced DSC values in each training batch. We kept this process stochastic to let the model see as many samples as possible without skewing its exposure to either low DSC or high DSC samples. This means we randomly sampled $n_{s}$ segmentations for bins that contained more than $n_{s}$ samples at each training iteration. Conversely, during the evaluation phase, we opted for a deterministic resampling approach, which enabled unbiased evaluations across various quality levels. In order to account for the variability introduced by the resampling process, we implemented a three-fold cross-validation strategy for all our experiments (see Section 3).

### v2

@@ 1 Introduction
Medical image segmentation plays an indispensable role for accurate diagnosis, monitoring, treatment planning, and population studies of diseases in modern medicine by enabling precise delineation of anatomical structures and pathological regions [18, 35]. In particular, precise segmentation of healthy and abnormal anatomy into multiple classes using magnetic resonance imaging (MRI) is crucial in this process. Recently, deep learning-based methods have achieved state-of-the-art performance in automated segmentation tasks including brain tumor [49, 26, 22, 23, 6] and cardiac [55, 27, 63] MRI segmentation. However, deep neural networks are sensitive to data distribution. This makes them prone to reduced performance when applied on out-of-distribution MRI scans due to variations in acquisition protocols, contrast, image quality, etc. Therefore, quality control (QC) is necessary to thoroughly assess segmentations before they are used for clinical purposes or large-scale research studies. QC tools are required to detect severe segmentation failures on a per-case basis, pinpoint areas needing segmentation refinement at the voxel level, and provide a quality measure for downstream analyses. Previously proposed QC methods fall into four main categories: uncertainty estimation-based, generative model-based, reverse classification accuracy (RCA)-based, and regression-based.
@@
The first category of automated segmentation QC methods operates on the premise that high uncertainty is reflective of poor quality segmentations [41, 3, 51, 40, 5, 50, 24, 39]. Accordingly, most studies focused on developing uncertainty measures to aggregate voxel-wise uncertainty as a proxy for segmentation quality measures, such as Dice Similarity Coefficient (DSC). However, most developed proxy measures [41, 50, 24] have not demonstrated a strong correlation with DSC. Moreover, non-negligible errors in uncertainty estimation were observed at the voxel level resulting in unreliable subject-level uncertainty aggregation [24]. Although there are also segmentation methods with built-in quality control [25], these methods can only be applied to deep learning-based models and thus cannot be used to assess the quality of segmentations obtained by other methods.
@@
The second category of automated QC methods is based on the assumption that there is a relationship between image intensities and tissue labels. Along that direction,  [19] proposed a Gaussian mixture model with hand-crafted features (i.e., geometric features, intensity features, gradient features, and ratio features) to characterize the segmentation quality and detect segmentation failure.  [61] proposed a variational autoencoder (VAE) to learn the latent representation of pairs of images and their ground-truth segmentations. During inference, the encoder is frozen while the decoder is refined for a given image-segmentation pair to produce a surrogate segmentation. The subject-level DSC is then computed between the query segmentation and the surrogate segmentation. The tuning of the decoder is required for each query image, which may be computationally expensive and time-consuming. Leveraging the advances in the image-to-image translation task, [31] proposed a generative adversarial network (GAN) to generate an informative reference image conditioned on the query segmentation mask that needs to be assessed. An auxiliary network (i.e., difference investigator) is then trained to predict image-level and pixel-level quality by taking the raw input image and the generated reference image as inputs. These approaches have only been validated in cardiac MRI segmentation QC that involves only a single modality and segmentation with regular shapes. Translating it to the more complex brain tumor segmentation QC scenario is difficult due to the presence of multiple modalities and intratumoral tissue heterogeneities, which are challenging to model.
Table 1: A detailed comparison of recent segmentation QC methods chosen from four different categories, including UE-based, generative-model-based, RCA-based, and regression-based methods. We chose recent representative methods from each category.
Tissue-specific error localization Support multiple quality measures Evaluated segmentation task Cross validation External evaluation Code availability
[24] coarse brain tumor
[61] cardiac
[31] cardiac
[57] whole body
[46] cardiac
[45] cardiac
[15] brain tumor
[29] brain tumor
[43] brain tumor
Ours brain tumor & cardiac
@@
The third category of the automated segmentation QC methods is built upon the reverse classification accuracy (RCA) framework [57]. This was initially designed for whole-body multi-organ MRI segmentation QC and was later applied to cardiac MRI datasets [46]. The RCA framework involves: (i) choosing a reference dataset with known ground-truth segmentation, (ii) training a segmenter using a query image-segmentation pair, (iii) using the trained segmenter to segment the images in the reference dataset, and (iv) estimating DSC as the maximum DSC achieved by the trained segmenter in the reference dataset. Although effective for whole-body multi-organ MRI segmentation and cardiac MRI segmentation QC, RCA’s reliance on a representative reference dataset and image registration poses challenges in brain tumors. The large variability in brain tumors makes the reference dataset hard to be representative. Different from whole-body multi-organ and cardiac segmentation, which have consistent shapes and appearances, the heterogeneous appearance and phenotypes of brain tumors complicate the establishment of correspondences for the tumorous areas between different subjects. Lastly, despite predicting subject-level DSC, the RCA framework lacks segmentation error localization at the voxel level.
@@
The fourth category of the segmentation QC methods is regression-based methods that directly predict the subject-level DSC. Early attempts employed a Support Vector Machine regression in combination with hand-crafted features to detect cardiac MRI segmentation failures [30, 3]. Instead of using hand-crafted features,  [45] proposed a convolutional neural network (CNN) regressor to automatically extract features from a large cardiac MRI segmentation dataset to predict DSC. Besides predicting DSC, [29] proposed a holistic rating to approximate how expert neuroradiologists classify high-quality and poor-quality segmentations. However, manually deriving holistic ratings by clinical experts is laborious and prone to inter-rater variability, hindering its applicability in large-scale datasets. There are also multi-dimensional regression QC approaches [see e.g., 15] that estimate segmentation quality measures for individual 2D slices and combine them to produce a 3D prediction. However, this method is specifically designed to predict the DSC, as its formulation relies on combining 2D DSC values into their 3D counterpart. Additionally, despite the fact that satisfactory performance was achieved in predicting subject-level segmentation metrics by these regression-based methods, voxel-level localization of segmentation error is unavailable.
@@
Although numerous efforts have been devoted to automated segmentation QC, the existing segmentation QC methods are still limited in many aspects (see a detailed comparison in Table 1). First, most studies have focused on evaluating on a single segmentation task, e.g., whole-body multi-organ MRI segmentation and cardiac MRI segmentation QC. Additionally, most of these approaches have not assessed out-of-distribution generalization on external datasets. Second, there has been limited research on segmentation QC specifically designed for brain tumor MRI segmentation QC. The heterogeneous and complex appearances of brain tumors, which vary in locations, sizes, and shapes, contribute to the increased challenge of QC in brain tumor MRI segmentation. Third, most prior QC approaches have focused on predicting subject-level DSC [24, 45, 46, 57, 15, 31] and do not consider or support the quality of the segmentation contour. However, both DSC and contour-based metrics (e.g., normalized surface dice (NSD)) are essential for a comprehensively assessment of the segmentation quality [38]. Fourth, it is crucial to address reliable voxel-level tissue-specific segmentation error localization. This localization is vital not only for auditing purposes but also for radiologists to prioritize cases that require manual refinement. By incorporating this component into quality control measures, it can enhance the accuracy and effectiveness of the overall segmentation process. However, this aspect has been largely overlooked in previous studies. One exception is the work proposed by [31]. However, this approach was only validated on cardiac segmentation QC with limited datasets. In addition, this method can only provide a binary segmentation error mask and is unable to identify voxel-level segmentation errors for different tissue classes, which limits its applicability in clinical practice.
@@
Fig. 1: Data generation for the brain tumor segmentation QC task: (a) The histogram of the DSC distribution in the generated BraTS training dataset before and after applying the resampling strategy. (b) Visual examples of the generated segmentation dataset ranging from low-quality to high-quality.
@@
To address these limitations, we proposed a novel deep learning model, termed QCResUNet, to jointly predict segmentation-quality metrics at the subject level and localize segmentation errors across different tissue classes at the voxel level. This study builds upon our previous preliminary work [43], but with a focus on developing a more comprehensive and applicable segmentation QC method. First, we extended the previous work to predict more comprehensive subject-level segmentation measures (DSC and NSD) as well as a collection of binary segmentation error maps, each corresponding to a different tissue class, to provide a more comprehensive evaluation of segmentation quality. To achieve this goal, we further proposed an attention-based segmentation error map aggregation mechanism to better delineate segmentation errors for different tissue classes. Second, we conducted rigorous evaluations to validate the generalizability of the proposed QCResUNet by performing three-fold cross-validation on the internal dataset and evaluating the proposed method on out-of-distribution datasets with MRI scans of varying image quality and segmentations produced. Third, we further examined the generalizability of the proposed QCResUNet by evaluating its ability to assess the segmentation quality for cardiac MRI. Fourth, we extensively evaluated the performance of the proposed method in comparison with several state-of-the-art segmentation QC methods, including RCA-based [57] and uncertainty-based methods [24]. Fifth, we performed an in-depth explainability analysis to elucidate the reasons behind the proposed method’s superior performance compared to other competing approaches.
The main contributions of this work are fivefold:
@@
  5. 5.
    To promote reproducibility and foster further research in the community, we have released our code at https://github.com/sotiraslab/QCResUNet.
@@ 2.1 Dataset
In this study, we used three datasets for evaluation of the QC performance on brain tumor segmentation. First, we used pre-operative multimodal MRI scans with gliomas of all grades (WHO Central Nervous System grades 2-4) grades from the BraTS 2021 challenge training dataset ($n=1251$). The BraTS dataset is a heterogeneous dataset consisting of cases from 23 different sites with various levels of quality and protocols. The BraTS dataset was used for training, validation, and internal testing. Additionally, we used two datasets that are not included in the BraTS 2021 dataset (i.e., the BraTS-SSA dataset and the WUSM dataset) for external testing, allowing for an unbiased assessment of the generalizability of the proposed method. The BraTS-SSA dataset ($n=40$[1] is an extension of the original BraTS 2021 dataset with patients from Sub-Saharan Africa, which includes lower-quality MRI scans (e.g., poor image contrast and resolution) as well as unique characteristics of gliomas (i.e., suspected higher rates of gliomatosis cerebri). The WUSM dataset ($n=175$) was obtained from the retrospective health records of the Washington University School of Medicine (WUSM), with a waiver of consent, in accordance with the Health Insurance Portability and Accountability Act, as approved by the Institutional Review Board (IRB) of WUSM (IRB no. PA18-1113). Each subject in all datasets comprised four modalities viz. pre-contrast T1-weighted (T1), T2-weighted (T2), post-contrast T1-contrast (T1c), and Fluid attenuated inversion recovery (FLAIR). In addition, multi-class tumor segmentation masks annotated by experts were also available. Segmentation masks delineated enhancing tumor (ET), non-enhancing tumor core (NCR), and edema (ED) classes. Following standard BraTS procedures, we combined the binary ET, NCR and ED segmentation masks to delineate the whole tumor (WT), tumor core (TC), and enhancing tumor. The WT mask consists of all tumor tissue classes (i.e., ET, NCR, and ED), while the TC mask comprises ET and NCR tissue classes.
@@
Scans from the BraTS training and BraTS-SSA datasets were already registered to the SRI24 anatomical atlas [48], resampled to 1-mm3 isotropic resolution and skull-stripped. For consistency, raw MRI scans from WUSM were pre-processed following the same protocol using the Integrative Imaging Informatics for Cancer Research: Workflow Automation for Neuro-oncology [11] framework. Subsequently, we z-scored all the skull-stripped scans in the BraTS datasets and the WUSM dataset on a per-scan basis. Finally, scans from the entire dataset were cropped to exclude background regions and then were zero-padded to a common dimension of $160\times 192\times 160$ using the nnUNet preprocessing pipeline [23].
@@
In evaluating the QC performance on the cardiac segmentation task, we used the Automated Cardiac Diagnosis Challenge (ACDC) dataset [8], which consists of 100 subjects (200 volumes). Each volume in the ACDC dataset is associated with a multi-class segmentation mask delineating the left ventricle (LV), myocardium (Myo), and right ventricle (RV). Lastly, each volume was cropped and zero-padded to a common dimension of $16\times 16\times 160$ using the nnUNet pipeline [23].
@@ 2.2 Segmentation Dataset Generation
As far as the brain tumor segmentation QC task is concerned, we first employed both CNN-based (i.e., nnUNet [23]) and Transformer-based (i.e., nnFormer [63]) segmentation UNets and trained them independently seven times by taking different modalities as input (namely T1-only, T1c-only, T2-only, FLAIR-only, T1-FLAIR, T1c-T2, and all four modalities). Since each modality is sensitive to certain tissue types (e.g., T1c is effective in detecting enhancing tumors and FLAIR is advantageous in identifying edema), this approach enabled us to produce segmentations with varying levels of quality. We selected these seven different combinations, as these generated sufficiently diverse segmentation samples. Empirical studies [see e.g., 11] have shown that including three input modalities does not significantly alter performance compared to only using two. Additionally, when including either T1c and FLAIR, the segmentation model can perform quite well, with the model that takes as input T1c and FLAIR performing as well as the model using all modalities. Based on these observations, we selected modality combinations that maximize segmentation quality diversity while discarding combinations that do not contribute additional variation, thereby reducing unnecessary computational burden.
Second, we sampled the segmentation along the training routines at various iterations to generate segmentation results from both fully trained and inadequately trained models. For this purpose, a small learning rate ($1~{}\times~{}10^{-6}$) was used during training to slower their convergence, allowing the segmentation to improve gradually from low quality to high quality. Third, we employed a data augmentation strategy based on image transformations to produce diverse segmentation results, termed SegGen. A series of image transformations, including rotation, translation, scaling, and elastic deformation, were randomly applied to each ground-truth segmentation three times with a probability of $0.5$ (see Table 2). Fourth, apart from the segmentations produced by the aforementioned three methods, we also created out-of-sample segmentation samples from a different segmentation method to assess the generalizability of our proposed model in the testing phase. To accomplish this, we employed the DeepMedic framework [26] to train seven segmentation models using the same diverse input modalities as we did in training the nnUNet and nnFormer. After applying the aforementioned four steps, we were able to create a set of erroneous segmentations for the cases included in the internal and external datasets. The erroneous segmentations along with the ground truth segmentations were subsequently used to estimate the DSC, NSD, and the ground truth segmentation error map (SEM) that were used to train the proposed network. The SEM was computed as the difference between a query segmentation (i.e., one of the erroneous segmentations we produced) and the corresponding ground truth. Examples of the generated segmentations can be found in Figure 1(b).
@@
Fig. 2: (a) The proposed QCResUNet model is a U-shaped neural network that takes as input $m$ imaging modalities ($\{M_{1},M_{2},\cdots,M_{m}\}$) and a multi-class segmentation mask to be evaluated ($S_{query}$). It generates three outputs: the subject-level segmentation-quality metrics DSC ($\text{DSC}_{pred}$) and NSD ($\text{NSD}_{pred}$) as well as a collection of $C$ binary voxel-level segmentation error maps $\{\text{SEM}_{\text{tissue}_{c}}\}_{c=1}^{C}$, each for each tissue class. In the case of brain tumor segmentation task, which is demonstrated as an example in this figure, QCResUNet takes as input four imaging modalities ($\{M_{1},M_{2},M_{3},M_{4}\}$) and a query segmentation $S_{query}$ that delineates WT, TC, and ET. It produces subject-level DSC and NSD, along with three binary SEMs corresponding to the segmentation error masks for WT, TC, and ET tissue classes. Note that the depicted image sizes as well as the number and size of convolution filters are specific to the brain tumor segmentation tasks. Subpanel figures (b) (c), and (d) depict the Residual Block employed in the encoder of QCResUNet, the Convolutional Block used in its decoder, and the Efficient Channel Attention (ECA) used for multiclass SEM aggregation, respectively. The abbreviations in the figure: Conv3D = 3D convolutional layer, GAP = global average pooling, FC = fully connected layer, LeakyReLU = leaky rectified linear unit, and One-hot Encoding = one-hot encodes the multi-class $S_{query}$ to a collection of binary masks.
@@
Despite our efforts to sample across different convergence stages by adopting a small learning rate, we observed that the deep learning models could successfully segment the majority of the cases. As a consequence, the resulting dataset was distributed unevenly across the different levels of quality, comprising mostly of higher quality segmentation results (see Figure 1(a)). Following the resampling strategy in [45], we randomly selected a subset of samples from each bin of the DSC histogram. This ensured that the number of segmentations from each bin was equal to the lowest count-per-bin value ($n_{s}$) across the distribution for all the segmentation datasets (refer to Figure 1(a)). For this purpose, we divided DSC values into 10 evenly spanned bins ranging from 0 to 1. We sampled cases from these bins such that we ensured that all bins contained an equal number of cases. For the training set, we did not manually select the DSC samples. Instead, at each training iteration, we randomly sampled segmentations that were evenly distributed across 10 bins to avoid imbalanced DSC values in each training batch. We kept this process stochastic to let the model see as many samples as possible without skewing its exposure to either low DSC or high DSC samples. This means we randomly sampled $n_{s}$ segmentations for bins that contained more than $n_{s}$ samples at each training iteration. Conversely, during the evaluation phase, we opted for a deterministic resampling approach, which enabled unbiased evaluations across various quality levels. In order to account for the variability introduced by the resampling process, we implemented a three-fold cross-validation strategy for all our experiments (see Section 3).